Question 7 Find the general solution of the given differential equation. y +2y +5y=0
Answer #1

yltzyt5y=0 (0²72D+5) Y=0 Auxillary Equation is Find m²+2m 1520 wots of Auxillay equation m² 72 m +5=0 mt2m +1+4=0 (m+1=-4 oh